The diagonal AC & BD of a rhombus intersect at (5,6). If A=(3,2). Then find the equation of diagonal BD.

Grade:11

1 Answers

Indu
7 years ago
A(3,2) and diagonals intersect at (5,6)
Since diagonals of a rhombus are perpendicular
Let m1 be the slope of AC
let m2 be the slope of BD
m1*m2=-1
m1=5-3/6-2=1/2
therefore; m2=-2
equation of BD is
(y-6)=m2(x-5)
(y-6)=-2(x-5)
2x+y+4=0